OpenStudy (anonymous):

y=x^2-4x+8 Find the vertex and x intercepts of the graph.

OpenStudy (luigi0210):

x-intercept: When y=0 Vertex: 2x-4 2x=4 x=2 The vertex is at x=2, so plug that in and find y

OpenStudy (anonymous):

Whats the formula to solve for y then? Would I just switch up the formula or...?

OpenStudy (cwrw238):

another way is to write it in the form (x - 2)^2 + 4

OpenStudy (luigi0210):

@cwrw238 Can finish explaining, I gtg

OpenStudy (anonymous):

So, is the formula to solve for y (x-2)^2+4?

OpenStudy (cwrw238):

this is called the vertex form of the equation the minimum value of (x - 2^2 is zero ( a square can't be negative) so at the vertex of the graph x = 2 and y = 4 vertex is at ( 2, 4)

OpenStudy (cwrw238):

the expression has a positive x^2 so will have a minimum value and graph will look like: 

OpenStudy (cwrw238):

there are no intercepts on x-axis if you try to solve (x - 2)^2 + 4 = 0 what do you get? (x - 2)^2 = = -4 can u find any real roots?

OpenStudy (anonymous):

no?

OpenStudy (cwrw238):

no - because there are no real roots of a negative number

OpenStudy (anonymous):

Okay thank you so much!

OpenStudy (cwrw238):

yw - have you any more questions about this problem?

OpenStudy (anonymous):

Nope